If f(x)=sinx then find (fofof)(a) where a belongs to [0, pi/2]
and d/dx(fofof)^x |x=a

Grade:11

1 Answers

Harshit Singh
3 years ago
Dear Student

f(x)=sinx then
(fofof)(a) = sin(sin(sin(a)))

d/dx(fofof)^x |x=a= cos(sin(sin a))* cos(sin a) * cos a

Thanks